for F(x) = 3x to the 2 power, find F(x+h)-f(x)
h
over h

Answer by nerdybill(7384)

for F(x) = 3x to the 2 power,
find [F(x+h)-f(x)]/h
.
f(x) = 3x^2 (given)
.
f(x+h) = 3(x+h)^2
f(x+h) = 3(x+h)(x+h)
f(x+h) = 3(x^2+2hx+h^2)
f(x+h) = 3x^2+6hx+3h^2
.
so, we substitute the above into:
[F(x+h)-f(x)]/h
.
[3x^2+6hx+3h^2 - 3x^2]/h
[6hx+3h^2]/h
6x+3h (this is what they're looking for)